# Caging in Time: A Framework for Robust Object Manipulation under Uncertainties and Limited Robot Perception

### v3

@@ Abstract
Real-world object manipulation has been commonly challenged by physical uncertainties and perception limitations. Being an effective strategy, while caging configuration-based manipulation frameworks have successfully provided robust solutions, they are not broadly applicable due to their strict requirements for the availability of multiple robots, widely distributed contacts, or specific geometries of the robots or the objects. To this end, this work proposes a novel concept, termed Caging in Time, to allow caging configurations to be formed even if there is just one robot engaged in a task. This novel concept can be explained by an insight that even if a caging configuration is needed to constrain the motion of an object, only a small portion of the cage is actively manipulating at a time. As such, we can switch the configuration of the robot strategically so that by collapsing its configuration in time, we will see a cage formed and its necessary portion active whenever needed. We instantiate our Caging in Time theory on challenging quasi-static and dynamic manipulation tasks, showing that Caging in Time can be achieved in general state spaces including geometry-based and energy-based spaces. With extensive experiments, we show robust and accurate manipulation, in an open-loop manner, without requiring detailed knowledge of the object geometry or physical properties, nor real-time accurate feedback on the manipulation states. In addition to being an effective and robust open-loop manipulation solution, the proposed theory can be a supplementary strategy to other manipulation systems affected by uncertain or limited robot perception.
@@ 1 Introduction
Different from approaches that focus on contacts, caging configuration-based manipulation has provided a novel paradigm to significantly reduce perception requirements. More importantly, although it does not aim at accurate control, caging configurations can robustly work by fully ignoring the effect of physical uncertainties, such as seen in grasping, in-hand manipulation, and multi-robot coordination tasks (Rodriguez et al., 2012; Song et al., 2021; Bircher et al., 2021). Concretely, a caging configuration aims at completely constraining all possible configurations of the target object within a known region (cage). The target object is manipulated as the cage moves or deforms, while the configuration of the target object is guaranteed to follow the cage to complete the task. However, it has been a challenge to apply this idea to general manipulation tasks, since forming a caging configuration has very strict requirements on the hardware, such as multiple robots to coordinate, widely distributed contacts, or specific geometries of the robot or the object to allow such configurations to exist (Makita and Wan, 2018).
@@
Figure 2: The theory of Caging in Time visualized through an example planar pushing task. I: A virtual cage, formed by line-shaped bars (grey), robustly pushes an object through a circular trajectory. II: Along the time dimension, we can see that only one bar (red) is effectively making contacts at a time, while other bars seem to be unnecessary. III: If we collapse all configurations of the effective bars (red) through time, a cage is formed, which can be unrolled into time to achieve the task in I with only one red bar at a time.
@@
This work proposes a novel concept, termed Caging in Time, to extend caging configuration-based manipulation to more general problems without hardware-specific assumptions. Example applications are shown in Fig. 2. The high-level idea of this new theory can be explained as follows. In an extreme situation, let us assume we have an object to manipulate and a robot is able to make an infinite number of contacts everywhere on the object. As such, the object is fully caged, and arbitrary manipulation can be achieved by moving all contacts simultaneously. In another extreme situation, let us assume a robot can make one contact with the object at a time, but it is able to switch the contact to other locations infinitely fast so that virtually there are contacts everywhere on the object. Equivalently to the former case, arbitrary manipulation can also be achieved with this virtual cage. Our idea of Caging in Time exploits the possibilities between these two extremes: We assume a robot can make one or a few contacts at a time, and it can switch to other contacts fast enough as needed so that, in time, it makes a cage. This idea is visualized in Fig. 2 via a planar pushing task, where an object is pushed by a virtual cage (grey) with multiple bars through a circular trajectory. Note that, physically and in time, only one bar (red) is effectively needed at a time to complete the task. Furthermore, by collapsing the configurations of the effective bars through time, a cage is formed and can be unrolled in time to complete the task as if a complete cage has always been there.
@@
We instantiated the Caging in Time theory on both quasi-static and dynamic manipulation tasks with a real robot. It is worthwhile to note that, the example instantiations showed that Caging in Time can be applied on general state spaces, including geometry-based and energy-based spaces. Without any sensing feedback, Caging in Time showed guaranteed task success on all experimented tasks, even when the manipulation is physically affected by in-task perturbations and unknown object shape variations. In comparison with a baseline closed-loop control approach, we show that our framework is similarly accurate, while being significantly more robust against perception uncertainties, as enabled by zero reliance on precise sensing feedback. Contributions: The proposed Caging in Time concept makes three theoretical contributions: 1) providing a planning paradigm for robust robot manipulation that can entirely remove the effect of perception uncertainties; 2) broadening the traditionally narrowly scoped caging configuration-based manipulation to a more general manipulation framework as enabled by strategic sequential robot motions; and 3) offering a possibility to precisely control robot manipulation without requiring sensing feedback, for the first time, to support robust manipulation in scenarios where perception is not reliable, e.g., object tracking in real-world is typically noisy and can lose track once in a while due to occlusions. Limitations: The work reported in this paper is the first step towards general applications of the Caging in Time concept. With an emphasis on the derivation of the foundations of the theory, this work does not develop algorithms for addressing general manipulation problems via Caging in Time skills. Specifically, we identify the following major limitations of the current scope of this work and state them before the details of the proposed theory: 1) as example instantiations of the proposed theory, the reported manipulation algorithms were designed for implementing Caging in Time on specific tasks only; and 2) manipulation physics or dynamics models are analytically derived for the purpose of explicitly verifying the theory, although simulation or learning-based models can be more efficient. Nevertheless, we underscore that the proposed Caging in Time framework is a complete theory for general manipulation problems as discussed in Sec. 4.
@@ 2 Related Works
Robot Manipulation: Contact properties, object geometrical and physical properties, and perfect object perception, are typically assumed to be sufficiently provided for modeling robot manipulation systems (Suomalainen et al., 2022; Bütepage et al., 2019). As an alternative, learning-based approaches assume that the data used for training the manipulation models are sufficiently covering all potential variations in relevant task domains, and that the perception system used in the task is the same or similar to the ones used in the training set, e.g., similar camera poses (Lee et al., 2019; Kaelbling, 2020; Andrychowicz et al., 2020; Kroemer et al., 2021). Along the other line, interactive perception has significantly reduced the requirements on certain prior knowledge and direct perception in manipulation systems (Bohg et al., 2017; Hang et al., 2021), while still assuming reliable perception through some channels to be available for robots to iteratively estimate the state of the manipulation system. Although many manipulation approaches have shown excellent performance, there is always a real-world challenge of unreliable perception, such as noise in object tracking or object occlusions, that will significantly affect the robustness of most manipulation systems. Different from most existing works, this work aims to use caging configurations along the time dimension to eliminate the reliance on, hence the negative effect of, unreliable perception. Note that the goal of this work is not about replacing other planning or control methods for manipulation with Caging in Time. Rather, we aim to provide an alternative or supplementary approach alongside others to improve the robustness of robot manipulation under unreliable perception.
Caging Configuration-based Manipulation: Caging configuration was proposed originally as a tool for a team of robots in $SE(2)$ to constrain the motion of a target object, so that the motion of the robot team will determine the motion of the caged object (Pereira et al., 2004; Sudsang and Ponce, 2000). Such a concept was later extended into the problem of robot grasping in 2D, for which the goal was to fix the target object within a small region (Rodriguez et al., 2012). By further modeling the topological relationships between robots and objects, hooking or latching-based manipulation was developed based on caging in topological spaces (Stork et al., 2013). More recently, a specific type of planar dexterous manipulation has been enabled by in-hand caging configurations so that the object is manipulated through controlled energy states inside the cage that deform over time (Bircher et al., 2021). With a mobile robot actively trying to escape from a team of other robots, caging configurations have also enabled formation-based motion planning to herd such a robot to a goal region (Song et al., 2021). Essentially, all existing frameworks aim at first constructing a cage to enclose the target object, and then maintain the cage while moving or deforming it to achieve a manipulation task. Lately, caging configuration-based manipulation has been extended to incorporate robot motions to relax the requirements on specific geometric properties of robots or objects (Dong and Pokorny, 2024; Dong et al., 2024). This line of research is focused on designing metrics to predict the manipulation robustness, while autonomous caging planning and control are left as open questions.
Challenges with Traditional Caging: The most fundamental limitation of traditional caging is that multiple robots (ZhiDong and Kumar, 2002), widely distributed contacts (Rodriguez et al., 2012), or specific geometries of the robots or objects (Varava et al., 2016) are required, making caging a concept not applicable to more general task setups. In this work, with the proposed framework of Caging in Time, we show that even with just one robot, we will be able to virtually construct cages in time to satisfy the aforementioned requirements. Additionally, traditional methods require very complex algorithms to verify caging configurations (Varava et al., 2021), which can be even harder for partial cages (Makita and Nagata, 2015). As will be seen with Caging in Time, since we can, in theory, virtually have our robots or contacts everywhere as needed, caging verification is generally a much simpler problem as we only need to make sure that the object does not penetrate the predefined cage boundaries.
@@ 3.1 Traditional Caging as Object Closure
In the scenario of an object caged by multiple robots, we denote the configuration space (C-space) of an object by $\mathcal{C}_{obj}$ and the configuration of an object (e.g., position and orientation) by $c_{obj}\in\mathcal{C}_{obj}$. The C-space and the configuration of the $i$-th robot are denoted by $\mathcal{C}_{i}$ and $c_{i}\in\mathcal{C}_{i}$, respectively. The object and all the robots share a common workspace $\mathcal{W}$, where $\mathcal{A}_{obj},\mathcal{A}_{i}\subset\mathcal{W}$ are the geometries of the object and of the $i$-th robot in the workspace. As such, the free C-space of the object $\mathcal{C}_{free}\subset\mathcal{C}_{obj}$ can be defined as the set of all possible states where the object does not collide with any robot:
@@
An illustration of the notations above is given in Fig. 4.
The traditional definition of caging, initially termed object closure, was first introduced by ZhiDong and Kumar (2002) as a condition for an object to be trapped by robots. In other words, the caging condition is met when there is no viable path for the object to move from its current configuration to a configuration infinitely far away.
@@ 3.2 Generalized Representations
We denote the state space of an object by $\mathcal{S}_{obj}$, which encompasses a wider range of properties than $\mathcal{C}_{obj}$. The state of an object at time $t$ is represented by $\mathbf{q}_{t}\in\mathcal{S}_{obj}$. Depending on the specific manipulation task, $\mathbf{q}_{t}$ may include not only configuration (e.g., position and orientation) but also velocity, acceleration, or other relevant properties.
@@ 4 Caging in Time
In this section, we will formally introduce the definition of our proposed Caging in Time theory. With as few as only a single robot interacting with the object, Caging in Time verifies that the object’s state remains caged while being manipulated, thereby enabling open-loop manipulation and guaranteeing the desired movement of the manipulated object. As an essential component of our proposed theory, we need to predict the bounded motions of the object, which involves propagating the PSS over time, as will be introduced in Sec. 4.1. Then, we will formally define Caging in Time in Sec. 4.2.
@@ 4.1 Propagation of PSS
At time $t$, for each specific state of the object in the PSS $\mathbf{q}_{t}\in\mathcal{Q}_{t}$, we denote the set of all allowed motions that transit the object’s state by $\mathcal{V}_{\mathbf{q}_{t}}$. We define the notation $\mathrm{T}_{x}\mathcal{M}$ to represent the tangent space of a manifold $\mathcal{M}$ at a point $x\in\mathcal{M}$ in this manifold. Note that $\mathcal{V}_{\mathbf{q}_{t}}\subset\mathrm{T}_{\mathbf{q}_{t}}\mathcal{S}_{obj}$, i.e., $\mathcal{V}_{\mathbf{q}_{t}}$ lies in the tangent space of the object’s state space at the state $\mathbf{q}_{t}$. Furthermore, based on the tangent bundle theory (Jost, 2008), we define the motion bundle at time $t$ in Eq. (4), which is the set consisting of all possible state-motion pairs of the object:
@@
We need to define a propagation function $\pi:\mathcal{QV}_{t}\mapsto\mathcal{S}_{obj}$ to predict the object’s state $\mathbf{q}_{t+1}$ at the next time step given a possible state $\mathbf{q}_{t}$ and a motion $\mathbf{v}_{t}$ of the object, i.e., $\mathbf{q}_{t+1}=\pi(\mathbf{q}_{t},\mathbf{v}_{t})$.
Figure 5: An illustration of PSS propagation. Left: The propagation for a single state $\mathbf{q}_{t}\in\mathcal{Q}_{t}\subset\mathcal{S}_{obj}$. With a set of possible motions $\mathcal{V}_{\mathbf{q}_{t}}$, $\mathbf{q}_{t}$ can be propagated to a set of different states (the orange region). Right: The propagation from the entire PSS $\mathcal{Q}_{t}$ to $\mathcal{Q}_{t+1}$ by propagating all points in $\mathcal{Q}_{t}$. The different colors of regions in $\mathcal{Q}_{t+1}$ illustrate the propagation from multiple different $\mathbf{q}_{t}\in\mathcal{Q}_{t}$.

### v4

@@ Abstract
Real-world object manipulation has been commonly challenged by physical uncertainties and perception limitations. Being an effective strategy, while caging configuration-based manipulation frameworks have successfully provided robust solutions, they are not broadly applicable due to their strict requirements on the availability of multiple robots, widely distributed contacts, or specific geometries of robots or objects.
Building upon previous sensorless manipulation ideas and uncertainty handling approaches, this work proposes a novel framework termed Caging in Time to allow caging configurations to be formed even with one robot engaged in a task. This concept leverages the insight that while caging requires constraining the object’s motion, only part of the cage actively contacts the object at any moment. As such, by strategically switching the end-effector configuration and collapsing it in time, we form a cage with its necessary portion active whenever needed.
We instantiate our approach on challenging quasi-static and dynamic manipulation tasks, showing that Caging in Time can be achieved in general cage formulations including geometry-based and energy-based cages. With extensive experiments, we show robust and accurate manipulation, in an open-loop manner, without requiring detailed knowledge of the object geometry or physical properties, or real-time accurate feedback on the manipulation states. In addition to being an effective and robust open-loop manipulation solution, Caging in Time can be a supplementary strategy to other manipulation systems affected by uncertain or limited robot perception.
@@ 1 Introduction
Unlike approaches that focus on contacts, caging configuration-based manipulation has provided a novel paradigm to significantly reduce perception requirements. More importantly, although it does not aim at accurate control, caging configurations can robustly work by fully ignoring the effect of physical uncertainties, such as seen in grasping, in-hand manipulation, and multi-robot coordination tasks (Rodriguez et al., 2012; Song et al., 2021; Bircher et al., 2021). Concretely, a caging configuration aims at completely constraining all possible configurations of the target object within a known region (cage). The target object is manipulated as the cage moves or deforms, while the configuration of the target object is guaranteed to follow the cage to complete the task (Wang et al., 2005). However, it has been a challenge to apply this idea to general manipulation tasks, since forming a caging configuration has very strict requirements on the hardware, such as multi-agent coordination, widely distributed contacts, or specific geometries of the robot or the object to construct such configurations (Makita and Wan, 2018).
@@
This work proposes a novel concept, termed Caging in Time, to extend caging configuration-based manipulation to more general problems without hardware-specific assumptions. Example applications are shown in Fig. 1. The high-level idea of this framework can be explained as follows. In an extreme situation, let us assume we have an object to manipulate and a robot is able to make an infinite number of contacts everywhere on the object. As such, the object is fully caged, and arbitrary manipulation can be achieved by moving all contacts simultaneously. In another extreme situation, let us assume a robot can make one contact with the object at a time, but it is able to switch the contact to other locations infinitely fast so that virtually there are contacts everywhere on the object. Equivalently to the former case, arbitrary manipulation can also be achieved with this virtual cage. Our idea of Caging in Time exploits the possibilities between these two extremes: We assume a robot can make one or a few contacts at a time, and it can switch to other contacts fast enough as needed so that, in time, it makes a cage. This idea is visualized in Fig. 2 via a planar pushing task, where an object is pushed by a virtual cage (grey) with multiple bars through a circular trajectory. Note that, physically and in time, only one bar (red) is effectively needed at a time to complete the task. Furthermore, by collapsing the configurations of the effective bars through time, a cage is formed and can be unrolled in time to complete the task as if a complete cage has always been there.
Figure 2: The theory of Caging in Time visualized through an example planar pushing task. I: A virtual cage, formed by line-shaped bars (grey), robustly pushes an object through a circular path. II: Along the time dimension, only one bar (red) is effectively making contacts at a time, while other bars seem to be unnecessary. III: If we collapse all configurations of the effective bars (red) through time, a cage is formed, which can be unrolled into time to achieve the task in I with only one bar at a time.
@@
We instantiated the Caging in Time theory on both quasi-static and dynamic manipulation tasks with a real robot. It is worthwhile to note that, the example instantiations showed that Caging in Time can be applied on general cage formulations, including geometry-based and energy-based cages. Without any sensing feedback, Caging in Time showed guaranteed task success on all experimented tasks, even when the manipulation is physically affected by in-task perturbations and unknown object shape variations. In comparison with a baseline closed-loop control approach, we show that our framework is similarly accurate, while being significantly more robust against perception uncertainties, as enabled by zero reliance on precise sensing feedback.
Contributions: The proposed Caging in Time concept makes three contributions: 1) providing a planning paradigm for robust robot manipulation that can significantly mitigate the effect of perception uncertainties; 2) broadening the traditional caging configuration-based manipulation to a more general manipulation framework as enabled by strategic sequential robot motions; and 3) offering an option for manipulation without relying on sensing feedback to support robust manipulation in various real-world tasks, especially in scenarios where perception is not reliable.
Limitations: The work reported in this paper is the first step towards general applications of the Caging in Time concept. With an emphasis on the derivation of the foundations of the theory, this work does not develop algorithms for addressing general manipulation problems via Caging in Time skills. Specifically, we identify the following major limitations of the current scope of this work and state them before the details of the proposed theory: 1) as example instantiations of the proposed theory, the reported manipulation algorithms were designed for implementing Caging in Time on specific tasks only; 2) manipulation physics or dynamics models are analytically derived for the purpose of explicitly verifying the theory, although simulation or learning-based models can be more efficient; and 3) this work mainly handles perception uncertainties while uncertainties in action execution and certain environmental interactions are not considered. Nevertheless, we underscore that the proposed Caging in Time framework is a complete theory for general manipulation problems as discussed in Sec. 4 and Sec. 8.
@@ 2 Related Works
Perception Assumptions in Manipulation: Contact properties, geometric and physical properties of objects, and perfect perception are commonly presumed when modeling manipulation systems (Suomalainen et al., 2022; Bütepage et al., 2019). Alternatively, recent learning-based approaches instead assume training data adequately covers all relevant task variation domains, with perception systems matching those in training environments, e.g., similar camera poses (Lee et al., 2019; Kaelbling, 2020; Andrychowicz et al., 2020; Kroemer et al., 2021). Meanwhile, interactive perception has significantly reduced the requirements for certain prior knowledge and direct perception (Bohg et al., 2017; Hang et al., 2021), while still assuming reliable perception through some channels for iterative state estimation. Despite demonstrating great performance in the lab, the robustness of many manipulation approaches typically faces significant challenges from unreliable perception in the real world, such as tracking noise, signal latency, and occlusions.
In early works, uncertainty was discussed in object pushing through analytical modeling (Lynch and Mason, 1996, 1995; Lynch, 1999; Akella and Mason, 1992), enabling robust planar pushing despite uncertainties. While these approaches handled some uncertainties in object-surface interactions, they still required precise object geometry and sensor feedback. Meanwhile, sensorless manipulation (Erdmann and Mason, 1988; Goldberg, 1993; Akella et al., 1997) provided inspiration for addressing manipulation without any feedback, with applications in orienting planar parts without sensors (Akella and Mason, 1998; Bohringer et al., 2000). However, these approaches were specifically limited to parts orienting, rather than general manipulation tasks. Later works on motion cones (Chavan-Dafle et al., 2020) and in-hand manipulation (Bhatt et al., 2021; Holladay et al., 2015) extended these concepts beyond pushing.
Notably, a recent work in robust pushing shares a similar idea with our work through variance-constrained optimization based on belief dynamics, generating stable pushing trajectories also in open-loop (J. et al., 2025). While theoretically rigorous, this approach remains confined to quasi-static pushing scenarios, without addressing its potential applications in broader manipulation contexts.
Therefore, a comprehensive framework that can robustly handle uncertainty even without requiring feedback and apply across diverse manipulation tasks is valuable. Different from most existing works, Caging in Time aims to use caging configurations along the time dimension to eliminate the reliance on, hence the negative effect of, unreliable perception. Importantly, our work is not intended to replace existing planning or control methods, but rather to complement them as a supplementary approach to enhance manipulation robustness under unreliable perception.
Caging Configuration-based Manipulation: Caging configuration initially emerged as a concept for multi-robot systems in $SE(2)$ to constrain the motion of a target object, enabling robust object transportation (Pereira et al., 2004; Sudsang and Ponce, 2000; Wang et al., 2005), or even herding mobile agents (Song et al., 2021). This concept was subsequently extended to planar grasping, with the objective of restricting the pose of a target object within a confined region (Rimon and Blake, 1999; Rodriguez et al., 2012; Varava et al., 2016; Stork et al., ), sometimes even with partially observable object geometries (Zarubin et al., 2013), as well as hooking and latching techniques developed through topological analysis (Stork et al., 2013). Recently, in-hand manipulation has been enhanced through caging configurations where objects are manipulated via controlled states within deformable cages (Komiyama and Maeda, 2021; Bircher et al., 2021). Besides algorithmic advances, the caging concept has also inspired specialized end-effector designs that enhance manipulation through physical embodiment (Dong et al., 2025; Xu et al., 2024). However, the most fundamental limitation of traditional caging is that multiple robots (ZhiDong Wang and Kumar, 2002), widely distributed contacts (Rodriguez et al., 2012), or specific geometries of the robots or objects (Varava et al., 2016) are required, making caging a concept not applicable to more general setups.
Beyond complete geometric caging, extended formulations have been proposed including energy-bounded caging to incorporate external forces like gravity (Mahler et al., 2016, ), and partial caging to relax full enclosure requirements (Varava et al., 2019, 2016) — both demonstrating that complete caging becomes unnecessary with environmental force assistance, broadening the application range for caging. These advancements necessitated practical metrics for escape probability evaluation (Varava et al., 2019; Stork et al., ), with sampling-based algorithms enabling quality assessment through escape path clearance analysis (Varava et al., 2020a, b). Such quantification methods have evolved beyond traditional caging to evaluate diverse manipulation tasks including pushing, soft object interaction, and multi-object handling (Dong et al., 2024, 2023; Dong and Pokorny, 2024). Despite effectively extending traditional formulations, energy-based partial caging still demands specific robot-object geometry knowledge and maintains complete constraints at each discrete time point, preserving specific end-effector geometry requirements. While robustness prediction metrics have advanced for more general tasks, the fundamental challenges of autonomous caging-based planning and control remain largely unresolved.
Building upon unified geometric and energy-based caging principles, Caging in Time departs from the conventional understanding that a cage must be complete at every moment. Instead, we establish a paradigm where a cage can be considered complete when it achieves completeness across the space-time continuum, thus significantly relaxing hardware requirements and expanding the practical applications of caging-based manipulation. Additionally, traditional methods require very complex algorithms to verify caging configurations (Varava et al., 2021), which can be even harder for partial cages (Makita and Nagata, 2015). As will be seen with Caging in Time, since we can, in theory, virtually have our robots or contacts everywhere as needed, caging verification is generally simpler as we only need to make sure that the object does not penetrate the predefined cage boundaries.
Theories for Uncertainty Handling: Our approach draws inspiration from foundational theoretical frameworks that address uncertainties through diverse mathematical formulations across control theory, motion planning, and decision-making. Belief space planning transforms state estimation into probability distributions for decision-making under uncertainty (Platt Jr et al., 2010; Kurniawati et al., 2008). POMDPs formalize partial observability by optimizing over belief states (Kaelbling et al., 1998; Silver and Veness, 2010). Contraction theory provides stability guarantees through analysis of convergence properties (Lohmiller and Slotine, 1998; Manchester and Slotine, 2017), while reachability analysis enables formal safety verification by computing attainable states (Mitchell et al., 2005; Althoff, 2010). LQR trees combine optimal control with sampling-based planning for stabilizing controllers (Tedrake et al., 2010; Majumdar and Tedrake, 2017), and set-based control ensures invariance properties for worst-case scenarios (Rakovic et al., 2005).
While all these approaches can effectively address uncertainties in different ways, they still rely on the assumption that perception feedback or certain prior geometric knowledge of the tasks is always, or at least partly, available. Inspired by these prior works and with an aim to address their limitations in contact-rich manipulation tasks, our proposed Caging in Time synthesizes their insights, including those in state representations and uncertainty-aware planning, from these fundamental theories into a framework that focuses on large perception uncertainties, in order to enable robust manipulation skills with practical real-world implementations.
@@ 3.1 Traditional Caging as Object Closure
In the scenario of an object caged by multiple robots, we denote the configuration space (C-space) of an object by $\mathcal{C}_{obj}$ and the configuration of an object (e.g., position and orientation) by $c_{obj}\in\mathcal{C}_{obj}$. The C-space and the configuration of the $i$-th robot are denoted by $\mathcal{C}_{i}$ and $c_{i}\in\mathcal{C}_{i}$, respectively. The object and all the robots share a common workspace $\mathcal{W}$, where $\mathcal{A}_{obj},\mathcal{A}_{i}\subset\mathcal{W}$ are the geometries of the object and of the $i$-th robot in the workspace. As shown in Fig. 3, the free C-space of the object $\mathcal{C}_{free}\subset\mathcal{C}_{obj}$ can be defined as the set of all possible states where the object does not collide with any robot:
@@
The traditional definition of caging, initially termed object closure, was first introduced by (ZhiDong Wang and Kumar, 2002) as a condition for an object to be trapped by robots. In other words, the caging condition is met when there is no viable path for the object to move from its current configuration to a configuration infinitely far away.
@@ 3.2 Generalized Representations
We denote the state space of an object by $\mathcal{S}_{obj}$, which encompasses a wider range of properties than $\mathcal{C}_{obj}$. The state of an object at time $t$ is represented by $\mathbf{q}_{t}\in\mathcal{S}_{obj}$. Depending on the specific manipulation task, $\mathbf{q}_{t}$ may include not only the configuration (e.g., position and orientation) but also velocity, acceleration, or other relevant properties.
@@ 4 Caging in Time
In this section, we will formally introduce the definition of our proposed Caging in Time theory. With as few as only a single robot interacting with the object, Caging in Time verifies that the object’s state remains caged while being manipulated, thereby enabling open-loop manipulation and guaranteeing the desired movement of the object. As an essential component of our theory, we need to predict the bounded motions of the object, which involves propagating the PSS over time, as will be introduced in Sec. 4.1. Then, we will formally define Caging in Time in Sec. 4.2.
@@ 4.1 Propagation of PSS
At time $t$, for each specific state of the object in the PSS $\mathbf{q}_{t}\in\mathcal{Q}_{t}$, we denote the set of all allowed motions that transit the object’s state by $\mathcal{V}_{\mathbf{q}_{t}}$. We define the notation $\mathrm{T}_{x}\mathcal{M}$ to represent the tangent space of a manifold $\mathcal{M}$ at a point $x\in\mathcal{M}$ in this manifold. Note that $\mathcal{V}_{\mathbf{q}_{t}}\subset\mathrm{T}_{\mathbf{q}_{t}}\mathcal{S}_{obj}$, i.e., $\mathcal{V}_{\mathbf{q}_{t}}$ lies in the tangent space of the object’s state space at the state $\mathbf{q}_{t}$. Based on the tangent bundle theory (Jost, 2011), we define the motion bundle at time $t$ in Eq. (4), which is the set consisting of all possible state-motion pairs of the object:
@@
We need to define a propagation function $\pi:\mathcal{QV}_{t}\mapsto\mathcal{S}_{obj}$ to predict the object’s state $\mathbf{q}_{t+1}$ at the next time step given a possible state $\mathbf{q}_{t}$ and a motion $\mathbf{v}_{t}$, i.e., $\mathbf{q}_{t+1}=\pi(\mathbf{q}_{t},\mathbf{v}_{t})$.
Figure 5: An illustration of the PSS propagation. Left: The propagation for a single state $\mathbf{q}_{t}\in\mathcal{Q}_{t}\subset\mathcal{S}_{obj}$. With a set of possible motions $\mathcal{V}_{\mathbf{q}_{t}}$, $\mathbf{q}_{t}$ can be propagated to a set of different states (the orange region). Right: The propagation from the entire PSS $\mathcal{Q}_{t}$ to $\mathcal{Q}_{t+1}$ by propagating all points in $\mathcal{Q}_{t}$. The different colors of regions in $\mathcal{Q}_{t+1}$ illustrate the propagation from multiple different $\mathbf{q}_{t}\in\mathcal{Q}_{t}$.